OpenStudy (raffle_snaffle):

True or False a) a*(axb)=b*(axb)=0

OpenStudy (raffle_snaffle):

I say true

OpenStudy (anonymous):

a(ab) = b(ab) = 0 a^2b = ab^2 = 0 True if a and b = 0

OpenStudy (anonymous):

What's the difference between * and x ? What kind of operations are they?

OpenStudy (raffle_snaffle):

it's taking the dot product and combining it with the cross product.

OpenStudy (anonymous):

Oh this is about vectors -.-

OpenStudy (raffle_snaffle):

It's my fault I didn't use the equations available to me.. I was super lazy not to use them. lol